### v2

@@ 3.3 Factor Model
4 Empirical Application
4.1 Data
We collect 120 monthly variables for the US spanning the period January 1959 to December 2023. The data, which are obtained from the Federal Reserve Economic Data (FRED) and are available at https://fred.stlouisfed.org, cover a wide range of key macroeconomic variables that applied economists monitor regularly, such as different measures of output, prices, interest and exchange rates, and stock market performance. We provide a full list of the data and their transformations in order to achieve stationarity in Appendix B. Out of the 120 series, we further distinguish a subset of “variables of interest”, that is, key variables of interest which we will inspect very closely in order to evaluate how well the different models perform. These variables, along with their transformation codes, are described in Table 2 below.
@@ 4.2 Forecasting exercise
Starting with the VARs, we estimate models of three different sizes: Medium (the 19 variables listed in Table 2), Large (the variables in medium plus an additional 20), and X-large (all 120 series available), that is, we consider 19, 39 and 120-variable sets. All models have a lag length of $p=6$. For the factor models, we follow the same strategy. Namely, for each variable size and each sample, we first extract the first common factor, $\hat{f}_{1t}$. Next, for each variable in that same set and each forecast horizon, we run the factor-augment autoregression in (11) and forecast up to $12$ months ahead.
We proceed in a similar fashion with all TSLMs. More specifically, for each TSLM considered, starting with the first forecast date (December 1984), we feed the historical data to the pre-trained model (when working with the univariate TSLMs we supplied one time series at a time, while with the multivariate TSLMs we supplied all time series within the specific model size at once). Next, we query the model to generate a zero-shot forecast for each of the variables provided, up to 12 steps ahead. [^13] For all our experiments, we downloaded the open-sourced pre-trained models and run them locally on our computer, using an Nvidia A6000 GPU with 48GB of memory. The only exception was Time-GPT, where we instead queried directly their API, with a separate query for each forecast date-time series combination. We also include a benchmark approach which uses OLS forecasts from univariate AR(1) models.
4.3 Measuring Predictive Accuracy
@@
where $\underline{t}$ and $\overline{t}$ denote the start and end of the out-of-sample period, and where $e_{i,j,\tau+h}^{2}$ and $e_{bcmk,j,\tau+h}^{2}$ are the squared forecast errors of variable $j$ at time $\tau$ and forecast horizon $h$ associated with model $i$ ($i\in\left\{\text{BVAR v1},\text{BVAR v2},\text{Factor model}\right\}\cup{TSLM}{}s$) and the AR(1) model, respectively. The point forecasts used to compute the forecast errors are obtained by averaging over the draws from the various models’ $h$-step-ahead predictive densities. Values of $RMSFE_{ijh}$ below one suggest that model $i$ produces more accurate point forecasts than the AR(1) benchmark for variable $j$ and forecast horizon $h$.
@@ 5.1 Do some LLMs perform better than others?
Figure 1: Distribution of RMSFEs (relative to an AR benckmark) for the TSLMs. The evaluation sample is Jan 1985 to December 2019.
@@
Table 3: Median, Std. deviation, Min and Max RMSFE by TSLM model type and forecast horizon. The evaluation sample is Jan 1985 to December 2019.
Outliers are not depicted for scaling reasons, and instead we decided to separately report the maximum and minimum RMSFEs attained by each model in Table 3 (together with their median and standard deviation). Indeed, it does happen on several occasions that for some series the TSLMs produce forecasts that are patently unreasonable. These forecasts would be automatically disregarded by the user, and would call for a fine tuning of the models. This result highlights that while zero-shot TSLMs may produce on average reasonable forecasts, they should not be used mechanically and do require careful monitoring.
@@ 5.2 Differences in point forecast between LLMs and econometric models
Figure 2: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs. The evaluation sample is Jan 1985 to December 2019.
Focusing first on the econometric models, it is remarkable that both the BVAR with natural conjugate prior and the DFM are consistently showing a good performance relative to the AR model. This is a well known result, which has been obtained by several different studies using this dataset. Surprisingly to us, the BVAR with asymmetric priors performs slightly worse than the symmetric counterpart, a result that is at odds with previous findings.
@@
Table 4: Median, Std. deviation, Min and Max RMSFE by model type and forecast horizon. The evaluation sample is Jan 1985 to December 2019.
@@ 5.3 Are the differences statistically significant?
Results for forecast horizons of one and three steps ahead are displayed in Table 5. At these short forecast horizons BVARs and DFMs outperform the AR benchmark in many cases (12, 13, and 14 cases respectively for BVAR(v.1), BVAR(v.2), and Factor model). In these instances, the forecasting gains are both large and statistically significant. In the remaining few cases in which the econometric models do not outperform the benchmark, the forecasting losses are small and never statistically significant. The TSLMs outperform the AR in fewer instances (5 for Moirai and 6 for TimesFM), also with gains that are both large and statistically significant. In the remaining cases, the forecasting losses are not statistically significant. As we have seen before, there is a tendency for TSLMs to produce the occasional more extreme results, with larger losses than those observed for the econometrics models.
@@
Results for forecast horizons of 6 and 12 step-ahead are displayed in Table 6. At these horizons, for the 19 series in this medium dataset, improvements versus the benchmark are less clear-cut. In general, neither the econometric models nor the TSLMs systematically outperform the AR benchmark, even though it is worth noting that there are no instances in which they significantly under perform. Overall, Moirai is the model that more often produced better forecasts (in 13 cases at the 6-step ahead and 10 cases at the 12-step ahead) followed closely by the BVAR(v1) and DFM. For some series, there are significant gains. In particular the TSLMs produce good forecasts for Personal Consumption Expenditures (PCEPI) and Hourly Earnings (CES0600000008), with gains that are significant and larger than those obtained by the econometric models. The econometric models produce good forecasts of Unemployment rate (UNRATE) and Weekly hours (CES0600000007).
@@
$h=1$
BVAR(v.1) BVAR(v.2) Factor model Moirai Large TimesFM
PAYEMS 0.84 *** 0.83 *** 0.87 *** 0.83 *** 0.82 ***
INDPRO 0.90 *** 0.93 ** 0.97 0.95 ** 0.94 **
FEDFUNDS 1.21 0.97 0.87 ** 1.00 1.23
UNRATE 0.88 *** 0.84 *** 0.89 *** 1.04 1.11
RPI 0.98 0.98 0.99 1.00 1.04
DPCERA3M086SBEA 1.00 0.99 0.98 * 1.02 1.06
CMRMTSPLx 0.96 * 0.95 ** 0.98 1.01 1.04
CUMFNS 0.90 * 0.91 ** 0.95 1.09 1.48
CES0600000007 0.90 *** 0.91 *** 0.91 ** 0.93 * 0.94 **
HOUST 0.95 *** 0.95 ** 0.93 *** 1.00 0.92 ***
S&P 500 1.04 1.03 1.00 1.03 1.02
T1YFFM 1.19 1.18 1.06 1.04 1.13
T10YFFM 1.06 1.00 1.00 1.04 1.16
BAAFFM 1.02 0.96 0.94 1.01 1.28
EXUSUKx 1.02 1.03 1.01 1.02 1.01
WPSFD49207 0.97 1.00 1.03 1.01 1.01
PPICMM 1.00 1.01 0.99 1.03 1.04
PCEPI 0.94 *** 0.94 ** 0.98 0.99 0.95 *
CES0600000008 0.85 *** 0.85 *** 0.78 *** 0.77 *** 0.78 ***
$h=3$
BVAR(v.1) BVAR(v.2) Factor model Moirai Large TimesFM
PAYEMS 0.76 ** 0.78 *** 0.79 ** 0.77 *** 0.75 **
INDPRO 0.94 0.98 0.99 0.97 0.94
FEDFUNDS 1.24 1.04 0.96 0.91 1.16
UNRATE 0.74 ** 0.72 *** 0.73 *** 0.91 0.78 **
RPI 1.00 1.00 0.99 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 0.99 1.00
CMRMTSPLx 0.97 ** 0.98 0.98 1.00 1.01
CUMFNS 0.79 ** 0.85 ** 0.92 0.93 1.32
CES0600000007 0.72 *** 0.79 *** 0.79 *** 0.79 *** 0.83 ***
HOUST 0.99 1.02 0.93 ** 1.01 0.78 ***
S&P 500 1.02 1.02 1.01 1.00 1.03
T1YFFM 1.32 1.30 1.14 1.03 1.12
T10YFFM 1.13 1.08 1.04 0.96 1.13
BAAFFM 1.09 1.05 1.01 0.93 1.17
EXUSUKx 1.03 1.05 1.01 1.02 1.02
WPSFD49207 0.99 1.02 1.05 0.99 1.03
PPICMM 1.00 1.02 1.00 1.00 1.03
PCEPI 0.93 ** 0.95 0.97 0.89 *** 0.91 ***
CES0600000008 0.83 *** 0.83 *** 0.77 *** 0.74 *** 0.75 ***
Table 5: Forecast horizons 1 and 3 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is Jan 1985 to December 2019.
$h=6$
BVAR(v.1) BVAR(v.2) Factor model Moirai Large TimesFM
PAYEMS 0.99 * 0.99 * 0.99 0.99 ** 0.99 *
INDPRO 0.99 1.00 1.01 0.99 0.98
FEDFUNDS 1.08 1.04 0.95 0.93 1.21
UNRATE 0.92 ** 0.92 *** 0.90 ** 0.95 0.94 **
RPI 1.00 1.00 1.00 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.00 1.00
CMRMTSPLx 1.00 1.00 1.00 1.01 1.02
CUMFNS 0.87 * 0.90 * 1.01 0.92 1.14
CES0600000007 0.69 *** 0.76 *** 0.78 *** 0.76 *** 0.79 ***
HOUST 1.01 1.05 0.94 * 0.98 0.66 **
S&P 500 1.01 1.01 1.00 1.00 1.03
T1YFFM 1.29 1.35 1.02 1.07 1.20
T10YFFM 0.98 1.05 0.92 *** 0.97 1.14
BAAFFM 1.03 1.06 0.98 0.93 1.15
EXUSUKx 1.02 1.04 1.00 1.00 1.02
WPSFD49207 1.01 1.04 1.03 0.99 1.02
PPICMM 1.02 1.04 1.00 1.01 1.06
PCEPI 0.92 0.96 0.92 0.85 *** 0.85 ***
CES0600000008 0.85 *** 0.84 *** 0.78 *** 0.76 *** 0.76 ***
$h=12$
BVAR(v.1) BVAR(v.2) Factor model Moirai Large TimesFM
PAYEMS 1.00 0.99 1.00 1.00 0.99
INDPRO 1.01 1.01 1.01 1.01 0.99
FEDFUNDS 0.96 1.01 0.96 1.00 1.08
UNRATE 0.89 *** 0.88 *** 0.86 ** 0.96 0.89 *
RPI 1.00 1.00 1.00 1.00 1.00
DPCERA3M086SBEA 1.00 1.00 1.00 1.00 1.00
CMRMTSPLx 1.01 1.01 1.01 1.02 1.02
CUMFNS 0.91 0.93 1.06 0.92 1.02
CES0600000007 0.70 *** 0.75 *** 0.81 *** 0.76 *** 0.78 ***
HOUST 1.03 1.11 1.02 0.96 0.48 **
S&P 500 1.01 1.01 1.00 1.01 1.05
T1YFFM 1.25 1.50 1.01 1.11 1.25
T10YFFM 0.87 1.05 0.90 ** 1.03 1.15
BAAFFM 0.94 1.01 0.96 0.96 1.22
EXUSUKx 1.00 1.01 1.00 1.03 1.00
WPSFD49207 1.02 1.05 1.03 1.00 1.03
PPICMM 1.00 1.00 1.01 1.01 1.03
PCEPI 0.90 * 0.97 0.90 * 0.82 *** 0.83 ***
CES0600000008 0.88 *** 0.88 *** 0.81 *** 0.78 *** 0.80 ***
Table 6: Forecast horizons 6 and 12 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is Jan 1985 to December 2019.
5.4 Are the differences stable over time?
The results presented so far gave an overall picture of the accuracy of the various models across the different forecast horizons, focusing on the whole evaluation sample, July 1987 to December 2019, at once. In this section, we perform two additional exercises to shed lights on the stability of the results presented. In our first exercise we continue to focus on the pre-COVID evaluation sample, but compute and inspect the Cumulative Sum of Squared prediction Error Differences (CSSED). This is a measure first proposed by Welch and Goyal (2008), and which allow to zoom in and identify the degree to which the relative ranking of the various models may have changed over time. We define CSSED as follows:
$CSSED_{i,j,t}=\sum_{\tau=\underline{t}}^{t}\left(e_{bcmk,j,\tau}^{2}-e_{i,j,\tau}^{2}\right).$ (14)
Positive and rising values of this measure indicate that the point forecasts generated by model $i$ for variable $j$ are more accurate than those produced by the AR(1) benchmark model.
Results are shown in Figure 3 and Figure 4. It is evident that there is a discontinuity point corresponding with the inception of the Great Financial Crisis, and that generally (albeit not always) the econometric models seem to be more robust to this event. With the exception of Unemployment rate, BVARs and Factor models respond and adjust better compared to TSLMs. Also, whilst the overall pattern emerging from the two class of models is similar, for some variables, TSLMs produce some unreasonable forecasts right from the start of the forecasting sample. Examples include Real personal income (RPI), Capacity Utilization (CUMFNS), Real Manufacturing and Trade Industries Sales (CMRMTSPLx), Weekly hours (CES0600000007), and the BAA-Fed fund rate spread (BAAFMM).
At the 12-step ahead horizon one can see a similar picture. For example, forecasts from TimesFM suddenly start deteriorating around 2003, and never recover. The pattern followed by the forecasts of the Unemployment rate is also remarkable, as it starts deteriorating in the nineties, then further deteriorates around 2003 to partially recover in 2010. There are also a few examples in which TimesFM suddenly improves its forecasting performance. As the mechanism trough which these models produce forecasts is highly nonlinear, it is hard to pin down the reason behind such sudden improvement (or deterioration) in the forecasts.
Figure 3: Cumulative sum of squared error differences for selected variables, $h=1$. The evaluation sample is Jan 1985 to December 2019.
Figure 4: Cumulative sum of squared error differences for selected variables, $h=12$. The evaluation sample is Jan 1985 to December 2019.
Figure 5: Distribution of RMSFEs (relative to an AR benckmark) for the TSLMs. The evaluation sample is Jan 2020 to December 2022 (excluding the months from March to June 2020).
Our second exercise focuses on the post COVID period sample, Jan 2020 to December 2022 (to reduce the impact on the results of the extreme outliers right around the onset of the pandemic, we excluded from our analysis the forecasts made between March and June 2020). [^15] The choice of excluding from the sample the most problematic COVID related observation is based exclusively on simplicity. Still, there is evidence that simple approaches such as this might work well, see Schorfheide and Song (2021). One could also consider more sophisticated ways to deal with the COVID observations such as Carriero et al. (2022) or Lenza and Primiceri (2022). Figure 5 and Figure 6 replicate our pre-COVID analysis, again first comparing the forecast accuracy of the various TSLMs, and next stacking that against the BVARs and factor model. Interestingly enough, the results are comparable to what we shown in Figure 1 and Figure 2, with Moirai-large and TimesFM still standing out among all TSLMs, and with a performance broadly comparable with that of the factor model and the BVAR with symmetric conjugate priors. It’s worth pointing out that in this sample the econometric models tend to make at times large forecast errors, mainly during the second half of 2020. In the case of the factor model it happens especially for interest rates and spreads. Instead, the TSLMs seem to perform relatively better in this more recent sample. However, the usual caveat about TSLMs should be kept in mind, i.e. the fact that, as shown in Table 1, these models do include the pandemic period in their training sample (and a few of these models also include may of the macroeconomic variables we set out to forecasts), and this makers it harder to pin down to which extent this forecasting performance could have been achieved in real time.
Figure 6: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs. The evaluation sample is Jan 2020 to December 2022 (excluding the months from March to June 2020).

### v3

@@ 3.3 Factor Model
3.4 Bayesian Additive Regression Trees
VARs and factor models assume linearity between endogenous variables and their lags or lags of their linear combinations. Although these models have proven quite successful at forecasting macroeconomic time series, the linearity assumption could be overly restrictive and this could have a negative impact on forecast accuracy. To test this hypothesis, we include Bayesian Additive Regression Trees (BART) in the pool of candidate models that we consider. BART (Chipman et al., 2010) models the conditional mean of the regression model by summing over a large number of trees which are, by themselves, constrained through a regularization prior. The main intuition behind the approach is that while each of these simple trees may only be able to explain a small fraction of the variation in the response variable, summing over a large number of them can describe the data extremely well. Prüser (2019), Huber and Rossini (2022), Huber et al. (2023) and Clark et al. (2023) are recent examples of successful applications of BART with macroeconomic data.
For each series $i$, we implement BART by estimating the following model [^13] We separate $y_{i,t-h}$ from the other predictors to avoid the model not recognizing that the lagged depedent variable is likely to have a non trivial role in predicting $y_{i,t}$. Our modeling choice is inspired by the approach of Carvalho et al. (2024), who modified the standard BART to introduce a Minnesota-type shrinkage specification into the node splitting selection. We also consider the alternative model $y_{i,t}=f\left(X_{t-h}\right)+\varepsilon_{i,t}$ but we found it to perform significantly below our preferred specification.
$y_{i,t}=\alpha_{h}+\gamma_{h}(L)y_{i,t-h}+f\left(X_{-i,t-h}\right)+\varepsilon_{i,t}$ (13)
where $X_{-i,t-h}$ denotes all series at time $t-h$ other than $y_{i,t-h}$ and $g\left(X_{-i,t-h}\right)$ is a non-linear function obtained by summing over a larger number of regression trees,
$f\left(X_{-i,t-h}\right)=\sum_{j=1}^{M}g\left(X_{-i,t-h}|\mathcal{T}_{j},m_{j}\right)$ (14)
where $g\left(X_{-i,t-h}|\mathcal{T}_{j},m_{j}\right)$ identifies a single tree model with $\mathcal{T}_{j}$ denoting the tree structure and $m_{j}$ is the vector of terminal node parameters associated with $\mathcal{T}_{j}$. Following, Chipman et al. (2010), we set $M=250$ in our application, and specify a prior on the tree structure along the lines they suggested. The $h$-step ahead forecast is then computed as
$\hat{y}_{i,t+h}=\hat{\alpha}_{h}+\hat{\gamma}_{h}y_{i,t}+\widehat{f}\left(X_{-i,t}\right)$ (15)
As with factor models, this approach requires a different regression for each forecast horizon.
@@ 3.4 Bayesian Additive Regression Trees
4 Empirical Application
4.1 Data
We collect 120 monthly variables for the US spanning the period January 1959 to December 2023. The data, which are obtained from the Federal Reserve Economic Data (FRED) and are available at https://fred.stlouisfed.org, cover a wide range of key macroeconomic variables that applied economists monitor regularly, such as different measures of output, prices, interest and exchange rates, and stock market performance. We provide a full list of the data and their transformations in order to achieve stationarity in Appendix B. Out of the 120 series, we further distinguish a subset of “variables of interest”, that is, key variables of interest which we will inspect very closely in order to evaluate how well the different models perform. These variables, along with their transformation codes, are described in Table 2.
@@ 4.2 Forecasting Exercise
Starting with the VARs, we estimate models of three different sizes: Medium (the 19 variables listed in Table 2), Large (the variables in medium plus an additional 20), and X-large (all 120 series available), that is, we consider 19, 39 and 120-variable sets. All models have a lag length of $p=6$. For the factor models, we follow the same strategy. Namely, for each variable size and each sample, we first extract the first common factor, $\hat{f}_{1t}$. Next, for each variable in that same set and each forecast horizon, we run the factor-augment autoregression in (13) and forecast up to $12$ months ahead. Similarly, for BART we re-estimate the trees for each variable set and sample, as the number of series in $X_{-i,t-h}$ changes with each variable set.
We proceed in a similar fashion with all TSLMs. More specifically, for each TSLM considered, starting with the first forecast date (December 1984), we feed the historical data to the pretrained model (when working with the univariate TSLMs we supplied one time series at a time, while with the multivariate TSLMs we supplied all time series within the specific model size at once). Next, we query the model to generate a zero-shot forecast for each of the variables provided, up to 12 steps ahead. [^14] For all our experiments, we downloaded the open-sourced pretrained models and run them locally on our computer, using an Nvidia A6000 GPU with 48GB of memory. The only exception was Time-GPT, where we instead queried directly their API, with a separate query for each forecast date-time series combination. We also include a benchmark approach which uses OLS forecasts from univariate AR(1) models.
@@
where $\underline{t}$ and $\overline{t}$ denote the start and end of the out-of-sample period, and where $e_{i,j,\tau+h}^{2}$ and $e_{bcmk,j,\tau+h}^{2}$ are the squared forecast errors of variable $j$ at time $\tau$ and forecast horizon $h$ associated with model $i$ ($i\in\left\{\text{BVAR v1},\text{BVAR v2},\text{Factor model},\text{BART}\right\}\cup{TSLM}{}s$) and the AR(1) model, respectively. The point forecasts used to compute the forecast errors are obtained by averaging over the draws from the various models’ $h$-step-ahead predictive densities. Values of $RMSFE_{ijh}$ below one suggest that model $i$ produces more accurate point forecasts than the AR(1) benchmark for variable $j$ and forecast horizon $h$.
@@ 5.1 Do some LLMs Perform Better than Others?
Figure 2: Distribution of RMSFEs (relative to an AR benckmark) for the TSLMs. The evaluation sample is January 1985 to December 2019.
@@
Table 3: Median, Std. deviation, Min and Max RMSFE by TSLM model type and forecast horizon. The evaluation sample is January 1985 to December 2019.
Outliers are not depicted for scaling reasons, and instead we decided to separately report the maximum and minimum RMSFEs attained by each model in Table 3 (together with their median and standard deviation). Indeed, it does happen on several occasions that for some series the TSLMs produce forecasts that are patently unreasonable. These forecasts would be automatically disregarded by the user and would call for a fine-tuning of the models. This result highlights that while zero-shot TSLMs may produce on average reasonable forecasts, they should not be used mechanically and do require careful monitoring.
@@ 5.2 Differences in Point Forecast between LLMs and Econometric Models
Figure 3: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs. The evaluation sample is January 1985 to December 2019.
Focusing first on the econometric models, it is remarkable that both the BVAR with natural conjugate prior and the factor model are consistently showing a good performance relative to the AR model. This is a well known result, which has been obtained by several different studies using this dataset. Surprisingly to us, the BVAR with asymmetric priors performs slightly worse than the symmetric counterpart, a result that is at odds with previous findings.
@@
BART 1.011 0.131 0.757 1.457 1.023 0.154 0.670 1.905
Table 4: Median, Std. deviation, Min and Max RMSFE by model type and forecast horizon. The evaluation sample is January 1985 to December 2019.
@@ 5.3 Are the Differences Statistically Significant?
Results for forecast horizons of one and three steps ahead are displayed in Table 5. At these short forecast horizons BVARs and factor models outperform the AR benchmark in many cases (12, 13, and 14 cases respectively for BVAR(v.1), BVAR(v.2), and Factor model). In these instances, the forecasting gains are both large and statistically significant. In the remaining few cases in which the econometric models do not outperform the benchmark, the forecasting losses are small and never statistically significant. The TSLMs outperform the AR in fewer instances (5 for Moirai and 6 for TimesFM), also with gains that are both large and statistically significant. In the remaining cases, the forecasting losses are not statistically significant. As we have seen before, there is a tendency for TSLMs to produce the occasional more extreme results, with larger losses than those observed for the econometrics models.
$h=1$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.84 *** 0.83 *** 0.87 *** 0.97 0.83 *** 0.82 ***
INDPRO 0.90 *** 0.93 ** 0.97 1.02 0.95 ** 0.94 **
FEDFUNDS 1.21 0.97 0.87 ** 1.34 1.00 1.23
UNRATE 0.88 *** 0.84 *** 0.89 *** 0.94 1.04 1.11
RPI 0.98 0.98 0.99 1.05 1.00 1.04
DPCERA3M086SBEA 1.00 0.99 0.98 * 1.03 1.02 1.06
CMRMTSPLx 0.96 * 0.95 ** 0.98 1.01 1.01 1.04
CUMFNS 0.90 * 0.91 ** 0.95 0.96 1.09 1.48
CES0600000007 0.90 *** 0.91 *** 0.91 ** 1.05 0.93 * 0.94 **
HOUST 0.95 *** 0.95 ** 0.93 *** 1.01 1.00 0.92 ***
S&P 500 1.04 1.03 1.00 1.03 1.03 1.02
T1YFFM 1.19 1.18 1.06 1.28 1.04 1.13
T10YFFM 1.06 1.00 1.00 1.09 1.04 1.16
BAAFFM 1.02 0.96 0.94 1.07 1.01 1.28
EXUSUKx 1.02 1.03 1.01 1.05 1.02 1.01
WPSFD49207 0.97 1.00 1.03 1.03 1.01 1.01
PPICMM 1.00 1.01 0.99 1.02 1.03 1.04
PCEPI 0.94 *** 0.94 ** 0.98 1.02 0.99 0.95 *
CES0600000008 0.85 *** 0.85 *** 0.78 *** 0.91 *** 0.77 *** 0.78 ***
$h=3$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.76 ** 0.78 *** 0.79 ** 0.99 0.77 *** 0.75 **
INDPRO 0.94 0.98 0.99 0.98 0.97 0.94
FEDFUNDS 1.24 1.04 0.96 1.09 0.91 1.16
UNRATE 0.74 ** 0.72 *** 0.73 *** 0.98 0.91 0.78 **
RPI 1.00 1.00 0.99 1.02 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 0.99 1.00
CMRMTSPLx 0.97 ** 0.98 0.98 1.01 1.00 1.01
CUMFNS 0.79 ** 0.85 ** 0.92 1.07 0.93 1.32
CES0600000007 0.72 *** 0.79 *** 0.79 *** 0.80 *** 0.79 *** 0.83 ***
HOUST 0.99 1.02 0.93 ** 1.04 1.01 0.78 ***
S&P 500 1.02 1.02 1.01 1.03 1.00 1.03
T1YFFM 1.32 1.30 1.14 1.32 1.03 1.12
T10YFFM 1.13 1.08 1.04 1.05 0.96 1.13
BAAFFM 1.09 1.05 1.01 1.05 0.93 1.17
EXUSUKx 1.03 1.05 1.01 1.06 1.02 1.02
WPSFD49207 0.99 1.02 1.05 1.06 0.99 1.03
PPICMM 1.00 1.02 1.00 1.05 1.00 1.03
PCEPI 0.93 ** 0.95 0.97 0.98 0.89 *** 0.91 ***
CES0600000008 0.83 *** 0.83 *** 0.77 *** 0.91 *** 0.74 *** 0.75 ***
Table 5: Forecast horizons 1 and 3 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
$h=6$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 0.99 * 0.99 * 0.99 0.81 ** 0.99 ** 0.99 *
INDPRO 0.99 1.00 1.01 1.01 0.99 0.98
FEDFUNDS 1.08 1.04 0.95 1.36 0.93 1.21
UNRATE 0.92 ** 0.92 *** 0.90 ** 0.85 0.95 0.94 **
RPI 1.00 1.00 1.00 1.03 1.00 1.01
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.00 1.00 1.00 0.98 1.01 1.02
CUMFNS 0.87 * 0.90 * 1.01 1.01 0.92 1.14
CES0600000007 0.69 *** 0.76 *** 0.78 *** 0.87 *** 0.76 *** 0.79 ***
HOUST 1.01 1.05 0.94 * 0.99 0.98 0.66 **
S&P 500 1.01 1.01 1.00 1.05 1.00 1.03
T1YFFM 1.29 1.35 1.02 1.33 1.07 1.20
T10YFFM 0.98 1.05 0.92 *** 1.09 0.97 1.14
BAAFFM 1.03 1.06 0.98 1.07 0.93 1.15
EXUSUKx 1.02 1.04 1.00 1.09 1.00 1.02
WPSFD49207 1.01 1.04 1.03 1.07 0.99 1.02
PPICMM 1.02 1.04 1.00 1.04 1.01 1.06
PCEPI 0.92 0.96 0.92 1.00 0.85 *** 0.85 ***
CES0600000008 0.85 *** 0.84 *** 0.78 *** 0.83 *** 0.76 *** 0.76 ***
$h=12$
BVAR(v.1) BVAR(v.2) Factor model BART Moirai Large TimesFM
PAYEMS 1.00 0.99 1.00 1.00 1.00 0.99
INDPRO 1.01 1.01 1.01 1.04 1.01 0.99
FEDFUNDS 0.96 1.01 0.96 0.86 1.00 1.08
UNRATE 0.89 *** 0.88 *** 0.86 ** 0.99 0.96 0.89 *
RPI 1.00 1.00 1.00 1.01 1.00 1.00
DPCERA3M086SBEA 1.00 1.00 1.00 1.01 1.00 1.00
CMRMTSPLx 1.01 1.01 1.01 1.03 1.02 1.02
CUMFNS 0.91 0.93 1.06 1.08 0.92 1.02
CES0600000007 0.70 *** 0.75 *** 0.81 *** 0.80 *** 0.76 *** 0.78 ***
HOUST 1.03 1.11 1.02 1.18 0.96 0.48 **
S&P 500 1.01 1.01 1.00 1.04 1.01 1.05
T1YFFM 1.25 1.50 1.01 1.33 1.11 1.25
T10YFFM 0.87 1.05 0.90 ** 0.88 ** 1.03 1.15
BAAFFM 0.94 1.01 0.96 0.82 0.96 1.22
EXUSUKx 1.00 1.01 1.00 1.09 1.03 1.00
WPSFD49207 1.02 1.05 1.03 1.05 1.00 1.03
PPICMM 1.00 1.00 1.01 1.04 1.01 1.03
PCEPI 0.90 * 0.97 0.90 ** 0.94 0.82 *** 0.83 ***
CES0600000008 0.88 *** 0.88 *** 0.81 *** 0.91 ** 0.78 *** 0.80 ***
Table 6: Forecast horizons 6 and 12 step-ahead. RMSFE relative to the AR(1). Differences in accuracy that are statistically significant at 10%, 5%, and 1% levels are denoted by one, two, or three stars, respectively. The evaluation sample is January 1985 to December 2019.
@@
Results for forecast horizons of 6 and 12 step-ahead are displayed in Table 6. At these horizons, for the 19 series in this medium dataset, improvements versus the benchmark are less clear-cut. In general, neither the econometric models nor the TSLMs systematically outperform the AR benchmark, even though it is worth noting that there are no instances in which they significantly under perform. Overall, Moirai is the model that more often produced better forecasts (in 13 cases at the 6-step ahead and 10 cases at the 12-step ahead) followed closely by the BVAR(v1) and factor model. For some series, there are significant gains. In particular the TSLMs produce good forecasts for Personal Consumption Expenditures (PCEPI) and Hourly Earnings (CES0600000008), with gains that are significant and larger than those obtained by the econometric models. The econometric models produce good forecasts of Unemployment rate (UNRATE) and Weekly hours (CES0600000007).
@@
5.4 Results during the Covid-19 Pandemic
The results presented so far gave an overall picture of the accuracy of the various models across the different forecast horizons, focusing on the evaluation sample January 1985 to December 2019. In this section, we focus on the post Covid-19 period sample, Jan 2020 to December 2022 (to reduce the impact on the results of the extreme outliers right around the onset of the pandemic, we excluded from our analysis the forecasts made between March and June 2020). [^16] The choice of excluding from the sample the most problematic Covid-19 related observation is based exclusively on simplicity. Still, there is evidence that simple approaches such as this might work well, see Schorfheide and Song (2021). One could also consider more sophisticated ways to deal with the Covid-19 observations such as Carriero et al. (2022) or Lenza and Primiceri (2022).
Figure 4 and Figure 5 replicate our pre-Covid-19 analysis, again first comparing the forecast accuracy of the various TSLMs, and next stacking that against the BVARs and factor model. Compared to what we shown in Figure 2 and Figure 3 there is an overall deterioration in forecast accuracy for all the models, which all struggle in overperforming the benchmark. Within the LLMs, only Moirai-large stands out, with a performance comparable with that of the factor model and the BVAR with symmetric conjugate priors. It’s worth pointing out that in this sample the econometric models tend to make at times large forecast errors, mainly during the second half of 2020. In the case of the factor model it happens especially for interest rates and spreads.
Figure 4: Distribution of RMSFEs (relative to an AR benckmark) for the TSLMs. The evaluation sample is January 2020 to December 2022 (excluding the months from March to June 2020).
Figure 5: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs. The evaluation sample is January 2020 to December 2022 (excluding the months from March to June 2020).
The usual caveat about TSLMs should be kept in mind, i.e. the fact that, as shown in Table 1, these models do include the pandemic period in their training sample (and a few of these models also include may of the macroeconomic variables we set out to forecasts), and this makers it harder to pin down to which extent this forecasting performance could have been achieved in real time.